OpenStudy (anonymous):

How do you verify the trigonometric function d/dx (cotx)= -csc^2x

zepdrix (zepdrix):

Recall:\[\Large \cot x\quad=\quad\frac{\cos x}{\sin x}\]

zepdrix (zepdrix):

Quotient rule! :)\[\Large \frac{d}{dx}\left(\frac{\cos x}{\sin x}\right)\]

OpenStudy (anonymous):

Ok thanks